# MicroLad: 2D-to-3D Microstructure Reconstruction and Generation via Latent Diffusion and Score Distillation

### v3

@@ 1 Introduction
Predictive materials design relies on rigorous structure–property (SP) linkages—quantitative relationships that connect microstructural features (e.g., phase distributions, volume fractions, and morphologies) to bulk responses and their associated uncertainties [1, 2, 3]. To meet this need, Integrated Computational Materials Engineering (ICME) has significantly advanced, particularly through property simulations based on numerical methods such as the finite element method (FEM), finite difference method (FDM), finite volume method (FVM), and fast Fourier transform (FFT) [4, 5, 6, 7]. These simulations are often coupled with homogenization theories applied to representative volume elements (RVEs) of various microstructure types [8, 9], including fuel cell electrodes, battery electrodes, fiber-reinforced composites, metal matrix composites, and polycrystalline aggregates [10, 11, 12, 13, 14, 15]. However, a comprehensive understanding of SP linkages remains elusive due to the scarcity of representative microstructure data. Material microstructures exhibit substantial stochastic variability [1, 16], so examining only a restricted region—whether with 2D imaging techniques (SEM, EBSD, TEM [17, 18]) or 3-D methods such as X-ray [19, 20] and FIB/SEM [21]—offers only a narrow view of the possible morphologies and their associated properties.
@@
Descriptor-based microstructure characterization and reconstruction (MCR) addresses this data gap by generating statistically representative microstructures that match prescribed statistical measures (e.g., n-point correlation functions), using methods such as the Yeong–Torquato algorithm [22, 23] and more recent gradient-based variants [24, 25, 26, 27, 28]. The core idea is to generate new, statistically representative microstructures from limited data, thereby augmenting datasets for analysis and simulation. This enables researchers to (i) produce large ensembles of equivalent microstructures, (ii) explore hypothetical “interpolated” states between known morphologies, and (iii) reconstruct full 3D volumes from sparse 2D observations. A major advantage of descriptor-based MCR is its physical interpretability: the same descriptors used as loss terms can be directly compared with experimental measurements. However, the approach is limited by the information content of the chosen descriptors. For instance, relying solely on two-point correlations may not accurately characterize complex microstructures [25], as distinct microstructural patterns may share identical or nearly identical two-point statistics. Consequently, multiple physically distinct structures can serve as equally valid "solutions" to the descriptor-matching problem.
@@
Data-driven microstructure reconstruction approaches, by contrast, learn microstructure distributions directly from images. Widely used architectures include deep generative models such as variational autoencoders (VAEs) [29, 30, 31, 32, 33, 34] and generative adversarial networks (GANs) [35, 36, 37, 38, 39, 40]. For example, SliceGAN [36], a GAN-based framework, has demonstrated 2D-to-3D reconstruction for a variety of anisotropic and multiphase microstructures. Instead of manually defining descriptors, these models learn salient morphological features from the data and can synthesize new instances that remain statistically close to the training set. More recently, diffusion models have emerged as state-of-the-art generative frameworks, achieving remarkable performance across diverse tasks due to their high fidelity, diversity, and stable training dynamics [41, 42, 43, 44, 45, 46, 47]. These models operate by progressively corrupting and then recovering data through a predefined forward and reverse diffusion process, typically formulated as a Markovian process—as in denoising diffusion probabilistic models (DDPMs) [42]. In addition, latent diffusion models (LDMs) [48] perform the diffusion steps in a learned, lower-dimensional latent space, reducing the computational cost while preserving essential semantic and structural features. This allows them to generate samples that faithfully follow the underlying data manifold. Several studies have shown that 2D diffusion models can faithfully reproduce complex microstructures—spherical inclusions, ceramics, composites, grain structures, and more [49, 50]. Furthermore, to address the 2D-to-3D reconstruction challenge, Lee and Yun proposed a multi-plane denoising diffusion (MPDD) framework [51, 52]. Their approach denoises multiple orthogonal 2D planes simultaneously with a pretrained 2D diffusion model, enforcing spatial coherence across slices and thereby synthesizing statistically equivalent 3D microstructures.
@@
In this study, we introduce MicroLad, a latent diffusion framework that unifies 2D-to-3D microstructure reconstruction with inverse-controlled generation (Figure 1). Leveraging a pretrained VAE on 2D microstructure data, a 2D LDM is trained to generate 2D microstructures from latent variables. The framework then performs dimensionality expansion from 2D to 3D using multi-plane denoising diffusion within the latent space (L-MPDD), decoding spatially connected 2D latent variables into coherent 3D microstructure volumes. Furthermore, using the 3D volume reconstructed via L-MPDD as initialization, the framework enables inverse-controlled 2D-to-3D microstructure generation via score distillation sampling (SDS) [53, 54], which iteratively nudges each latent slice toward target microstructural descriptors and effective material properties.
@@
The remainder of the paper formalizes the forward and inverse problems in the 2D-to-3D dimensionality expansion setting, describes the MicroLad architecture, and presents quantitative results on reconstruction accuracy, inverse-controlled generation, and computational cost. We validate the framework on binary carbonate [55] and three-phase solid oxide fuel cell (SOFC) microstructures [56], demonstrating accurate 3D reconstruction and inverse-controlled generation. The results are then discussed in terms of reconstruction performance and the effectiveness of inverse-controlled generation.
@@ 2.1 Forward/inverse structure–property (SP) linkage
In conventional SP linkage approaches within ICME framework (Figure 2a), the workflow typically includes the following steps [1, 57]: 1) acquisition of experimental micrograph data of material microstructures, 2) micrograph data analysis, which frequently involves filtering and segmentation of microstructure images followed by statistical analysis to extract microstructural descriptors such as n-point correlation functions and volume fractions of material phases, and 3) analysis of material responses and effective properties based on the provided microstructures, commonly using numerical methods under defined boundary conditions. In this forward linkage, one can investigate how variations in microstructural features influence material properties, enabling the evaluation of property ranges and performance predictions under specific loading scenarios, often combined with uncertainty quantification techniques [58].
@@ 3.1 Denoising diffusion probabilistic models (DDPMs)
DDPM [42] defines a forward diffusion process as a Markov chain as follows:
@@ 3.2 Latent diffusion models (LDMs)
According to Lee and Yun [52, 51], MPDD (which performs iterative 2D denoising diffusion along the three Cartesian axes of a 3D volume) effectively approximates the 3D reverse diffusion process. In this study, we adopt a similar approach but implement MPDD within the latent space (i.e., L-MPDD) using a pretrained 2D LDM. Because the pretrained 2D LDM lacks an explicit mechanism for enforcing 3D connectivity (5), we leverage L-MPDD to implicitly achieve this goal by applying denoising diffusion along all three orthogonal latent dimensions: height $H^{\prime}$, width $W^{\prime}$, and depth $D^{\prime}$.
@@ 3.4 Inverse-controlled 2D-to-3D microstructure generation with score-distillation sampling
where $w_{M}$ and $w_{P}$ are weighting terms and $\eta$ is the learning rate. Here $\nabla^{\mathrm{pg}}$ denotes a pseudo-gradient update. In the SDS term, we treat $\varepsilon_{\theta}(\cdot)$ as a constant target and do not back-propagate through the frozen diffusion network (i.e., we drop the Jacobian $\partial\varepsilon_{\theta}/\partial\mathbf{z}_{\text{slice}}$). This mirrors the original SDS setup [53], where the diffusion prior is frozen and only the external generator is optimized. In our case, the “external generator” is the latent slice itself. The updated latent slice is decoded and replaces the original:
@@ 3.5 Implementation details
The VAE was pretrained on a variety of microstructure types—including Voronoi grains, generic grain microstructures, spherical inclusions, aligned elliptical inclusions, graphite, nickel-based superalloys (NBSA), Gaussian random fields (GRF), and random elliptical inclusions—using the open-source microstructure dataset from Robertson et al.[59]. It was then fine-tuned on the target microstructures of interest: binary carbonate and three-phase SOFC microstructures. Latent representations were encoded as $4\times 16\times 16$ feature maps. The LDM operates on this latent space and was trained for 1000 diffusion steps using a linear noise schedule $\beta\in[1\times 10^{-4},2\times 10^{-2}]$. Full architectural specifications and hyperparameters for both the VAE and LDM are provided in Appendix B.
@@ 3.6 Dataset preparation
For the binary microstructure case, patches were randomly sampled from a $318\times 337$ binary image of a carbonate structure [55] to demonstrate 2D-to-3D reconstruction using MicroLad. For the multiphase case, a SOFC anode microstructure [56] was selected, which consists of three phases: pores, nickel (Ni), and yttria-stabilized zirconia (YSZ). In SOFC anodes, each phase plays a distinct role in electrochemical performance: Ni serves as the electron-conducting phase and catalyzes hydrogen oxidation, YSZ is an oxygen-ion conductor that enables ion transport from the electrolyte, and the pore phase provides gas channels for fuel diffusion. A single $950\times 843$ microstructure image was taken from a reference SOFC dataset and manually segmented into the three constituent phases before being used for training. For each case, a total of 300 (binary) and 600 (three-phase) cropped images were randomly sampled, and 8-fold data augmentation was applied to enhance training diversity.
@@ 3.7 Metrics
Relative surface area is a commonly used microstructural descriptor that quantifies the interfacial complexity of each phase, which is particularly important in multiphase materials where phase boundaries influence functional performance (e.g., ion/electron transport pathways in SOFCs) [36, 56]. In our implementation, surface area is computed by applying Gaussian smoothing to the phase masks and computing their total variation (i.e., the gradient magnitude across the spatial domain) to enable differentiability. The surface area of each phase is then normalized to obtain its relative contribution.
@@
This formulation is sensitive to both phase connectivity and tortuosity [60, 61], making $D_{e}$ an indicator of microstructural functionality. Additional details on the FEM implementation are provided in Section C.
@@ 4 Results
To further validate the reconstruction performance of the proposed MicroLad framework, we compare its accuracy on three-phase SOFC microstructures with two existing methods: MPDD [52, 51] and SliceGAN [36]. As shown in Appendix D, the maximum relative error in terms of two-point correlation for MPDD is 12.44%, and for SliceGAN it is 11.01%. In contrast, MicroLad achieves lower errors when $S_{2}$ guidance is applied, demonstrating strong performance in 2D-to-3D microstructure reconstruction with respect to spatial correlation.
@@ 4.2 Inverse-controlled 2D-to-3D generated microstructure samples
Figure 6 presents the generated 3D microstructure volumes targeting the maximum relative surface area observed in the original dataset for each phase. In contrast to the $V_{f}$-guided generation, the surface area-guided results reveal a tendency for each material phase to form elongated, thin, and interconnected structures that maximize interfacial area with adjacent phases. This behavior also demonstrates that the proposed inverse-controlled generation framework can capture and reflect physically meaningful trends, as the surface area between phases plays a critical role in the electrochemical performance of SOFC anodes by influencing phase boundary reactions [56]. These results highlight MicroLad’s ability to control and analyze phase-specific morphological features and to generate realistic 3D structures suitable for studying interfacial complexity.
@@ 5.1 Distributional shifts and objective trends in latent space
To assess whether generation with the proposed MicroLad framework genuinely expands the microstructure analysis space, we can analyze the property distribution as done in Section 4.2. Additionally, we can examine the data distribution in the latent space of the LDM. To investigate the behavior of the inverse-controlled generation and the distance between the 2D-to-3D generated data (i.e., 2D slices within the 3D volume) and the original 2D data, we examine their latent embeddings using uniform manifold approximation and projection (UMAP) [62]. A detailed explanation of the latent embedding procedure and UMAP projection is provided in Appendix F.
@@
Figure 8(h) shows that the trend for increasing the effective diffusivity of the pore phase closely resembles those of $V_{f}^{\text{Pore}}$ and $SA^{\text{Pore}}$. This similarity is expected, as the volume fraction of the pore phase is relatively low compared to the YSZ and Ni phases, and the relative diffusivity of the pore phase in 2D slices is also quite small (with a maximum value around 0.0039). Consequently, achieving a higher $D_{e}^{\text{Pore}}$ effectively requires increasing both the pore volume fraction and surface area. In contrast, Figures 8(i) and 8(j) show more abrupt, threshold-like behavior for $D_{e}^{\text{YSZ}}$ and $D_{e}^{\text{Ni}}$, respectively. In these cases, the effective diffusivity remains low until a certain structural threshold is reached, beyond which the diffusivity rapidly increases. This is physically reasonable: once a given phase becomes sufficiently connected across the domain, forming a continuous pathway from one side to the other, its effective diffusivity improves significantly. This on/off behavior is particularly evident in $D_{e}^{\text{Ni}}$, where a cluster of high-diffusivity samples appears in the lower-left region of the plot in Figure 8(j). Although the diffusivity of YSZ remains lower than that of Ni—primarily due to its granular structure and presence of sintered necks [63, 56], as opposed to the more continuous network formed by Ni—a similar trend can still be observed in $D_{e}^{\text{YSZ}}$, as shown in Figure 8(i).
@@
To compare the diffusive behavior of each phase in the inverse-controlled generated 3D microstructures with those reconstructed (Recon.) without property guidance, the 3D tortuosity [64, 65] of both sets of volumes was computed along the three Cartesian axes (x, y, and z), as shown in Table 2. As observed, the reconstructed 3D volumes generally exhibit higher tortuosity (i.e., lower diffusivity), while the inverse-controlled samples show reduced tortuosity as the target $D_{e}$ is optimized for each material phase of interest. This observation suggests that enforcing target diffusivity at the slice level during 3D microstructure generation can effectively promote globally connected diffusive networks within the volume. Future work should explore more advanced control strategies, such as directionally weighted SDS loss or hierarchical optimization across multi-scale representations, to further enhance the spatial tailoring of 3D transport properties. More sophisticated control over 3D material behavior also warrants further investigation.

### v4

@@ 1 Introduction
Predictive materials design relies on rigorous structure–property (SP) linkages—quantitative relationships that connect microstructural features (e.g., phase distributions, volume fractions, and morphologies) to bulk responses and their associated uncertainties [bostanabad2018computational, shojaeefard2016review, bargmann2018generation]. To meet this need, Integrated Computational Materials Engineering (ICME) has significantly advanced, particularly through property simulations based on numerical methods such as the finite element method (FEM), finite difference method (FDM), finite volume method (FVM), and fast Fourier transform (FFT) [ghosh2023statistically, david2020application, chen2017finite, moulinec1998numerical]. These simulations are often coupled with homogenization theories applied to representative volume elements (RVEs) of various microstructure types [geers2010multi, kalamkarov2009asymptotic], including fuel cell electrodes, battery electrodes, fiber-reinforced composites, metal matrix composites, and polycrystalline aggregates [wargo2012selection, joos2012representative, pan2008numerical, jagadeesh2020review, kashkooli2017representative, benedetti2013three]. However, a comprehensive understanding of SP linkages remains elusive due to the scarcity of representative microstructure data. Material microstructures exhibit substantial stochastic variability [bostanabad2018computational, holm2020overview], so examining only a restricted region—whether with 2D imaging techniques (SEM, EBSD, TEM [zaefferer2011critical, brodusch2018imaging]) or 3-D methods such as X-ray [maire2014quantitative, stock2008recent] and FIB/SEM [xu2017enhanced]—offers only a narrow view of the possible morphologies and their associated properties.
@@
Descriptor-based microstructure characterization and reconstruction (MCR) addresses this data gap by generating statistically representative microstructures that match prescribed statistical measures (e.g., n-point correlation functions), using methods such as the Yeong–Torquato algorithm [torquato2002random, jiao2007modeling] and more recent gradient-based variants [seibert2022descriptor, seibert2022microstructure, ma2024stochastic, blumer2025generative, seibert2023two]. The core idea is to generate new, statistically representative microstructures from limited data, thereby augmenting datasets for analysis and simulation. This enables researchers to (i) produce large ensembles of equivalent microstructures, (ii) explore hypothetical “interpolated” states between known morphologies, and (iii) reconstruct full 3D volumes from sparse 2D observations. A major advantage of descriptor-based MCR is its physical interpretability: the same descriptors used as loss terms can be directly compared with experimental measurements. However, the approach is limited by the information content of the chosen descriptors. For instance, relying solely on two-point correlations may not accurately characterize complex microstructures [seibert2022microstructure], as distinct microstructural patterns may share identical or nearly identical two-point statistics. Consequently, multiple physically distinct structures can serve as equally valid "solutions" to the descriptor-matching problem.
@@
Data-driven microstructure reconstruction approaches, by contrast, learn microstructure distributions directly from images. Widely used architectures include deep generative models such as variational autoencoders (VAEs) [jung2021exploration, zhang2022reconstruction, ji2024towards, attari2023towards, noguchi2021stochastic, white2024exploring] and generative adversarial networks (GANs) [kench2022microlib, kench2021generating, hong2024mechanical, zheng2024generative, gayon2020pores, henkes2022three]. For example, SliceGAN [kench2021generating], a GAN-based framework, has demonstrated 2D-to-3D reconstruction for a variety of anisotropic and multiphase microstructures. Instead of manually defining descriptors, these models learn salient morphological features from the data and can synthesize new instances that remain statistically close to the training set. More recently, diffusion models have emerged as state-of-the-art generative frameworks, achieving remarkable performance across diverse tasks due to their high fidelity, diversity, and stable training dynamics [yang2023diffusion, ho2020denoising, dhariwal2021diffusion, song2020score, croitoru2023diffusion, cao2024survey, zhang2023text]. These models operate by progressively corrupting and then recovering data through a predefined forward and reverse diffusion process, typically formulated as a Markovian process—as in denoising diffusion probabilistic models (DDPMs) [ho2020denoising]. In addition, latent diffusion models (LDMs) [rombach2022high] perform the diffusion steps in a learned, lower-dimensional latent space, reducing the computational cost while preserving essential semantic and structural features. This allows them to generate samples that faithfully follow the underlying data manifold. Several studies have shown that 2D diffusion models can faithfully reproduce complex microstructures—spherical inclusions, ceramics, composites, grain structures, and more [dureth2023conditional, lee2024microstructure]. Furthermore, to address the 2D-to-3D reconstruction challenge, Lee and Yun proposed a multi-plane denoising diffusion (MPDD) framework [lee2024multi, lee2024denoising]. Their approach denoises multiple orthogonal 2D planes simultaneously with a pretrained 2D diffusion model, enforcing spatial coherence across slices and thereby synthesizing statistically equivalent 3D microstructures.
@@
In this study, we introduce MicroLad, a latent diffusion framework that unifies 2D-to-3D microstructure reconstruction with inverse-controlled generation (Figure 1). Leveraging a pretrained VAE on 2D microstructure data, a 2D LDM is trained to generate 2D microstructures from latent variables. The framework then performs dimensionality expansion from 2D to 3D using multi-plane denoising diffusion within the latent space (L-MPDD), decoding spatially connected 2D latent variables into coherent 3D microstructure volumes. Furthermore, using the 3D volume reconstructed via L-MPDD as initialization, the framework enables inverse-controlled 2D-to-3D microstructure generation via score distillation sampling (SDS) [poole2022dreamfusion, mcallister2024rethinking], which iteratively nudges each latent slice toward target microstructural descriptors and effective material properties.
@@
The remainder of the paper formalizes the forward and inverse problems in the 2D-to-3D dimensionality expansion setting, describes the MicroLad architecture, and presents quantitative results on reconstruction accuracy, inverse-controlled generation, and computational cost. We validate the framework on binary carbonate [prodanovic2015digital] and three-phase solid oxide fuel cell (SOFC) microstructures [hsu2018mesoscale], demonstrating accurate 3D reconstruction and inverse-controlled generation. The results are then discussed in terms of reconstruction performance and the effectiveness of inverse-controlled generation.
@@ 2.1 Forward/inverse structure–property (SP) linkage
In conventional SP linkage approaches within ICME framework (Figure 2a), the workflow typically includes the following steps [bostanabad2018computational, panchal2013key]: 1) acquisition of experimental micrograph data of material microstructures, 2) micrograph data analysis, which frequently involves filtering and segmentation of microstructure images followed by statistical analysis to extract microstructural descriptors such as n-point correlation functions and volume fractions of material phases, and 3) analysis of material responses and effective properties based on the provided microstructures, commonly using numerical methods under defined boundary conditions. In this forward linkage, one can investigate how variations in microstructural features influence material properties, enabling the evaluation of property ranges and performance predictions under specific loading scenarios, often combined with uncertainty quantification techniques [chernatynskiy2013uncertainty].
@@ 3.1 Denoising diffusion probabilistic models (DDPMs)
DDPM [ho2020denoising] defines a forward diffusion process as a Markov chain as follows:
@@ 3.2 Latent diffusion models (LDMs)
According to Lee and Yun [lee2024denoising, lee2024multi], MPDD (which performs iterative 2D denoising diffusion along the three Cartesian axes of a 3D volume) effectively approximates the 3D reverse diffusion process. In this study, we adopt a similar approach but implement MPDD within the latent space (i.e., L-MPDD) using a pretrained 2D LDM. Because the pretrained 2D LDM lacks an explicit mechanism for enforcing 3D connectivity (5), we leverage L-MPDD to implicitly achieve this goal by applying denoising diffusion along all three orthogonal latent dimensions: height $H^{\prime}$, width $W^{\prime}$, and depth $D^{\prime}$.
@@ 3.4 Inverse-controlled 2D-to-3D microstructure generation with score-distillation sampling
where $w_{M}$ and $w_{P}$ are weighting terms and $\eta$ is the learning rate. Here $\nabla^{\mathrm{pg}}$ denotes a pseudo-gradient update. In the SDS term, we treat $\varepsilon_{\theta}(\cdot)$ as a constant target and do not back-propagate through the frozen diffusion network (i.e., we drop the Jacobian $\partial\varepsilon_{\theta}/\partial\mathbf{z}_{\text{slice}}$). This mirrors the original SDS setup [poole2022dreamfusion], where the diffusion prior is frozen and only the external generator is optimized. In our case, the “external generator” is the latent slice itself. The updated latent slice is decoded and replaces the original:
@@ 3.5 Implementation details
The VAE was pretrained on a variety of microstructure types—including Voronoi grains, generic grain microstructures, spherical inclusions, aligned elliptical inclusions, graphite, nickel-based superalloys (NBSA), Gaussian random fields (GRF), and random elliptical inclusions—using the open-source microstructure dataset from Robertson et al.[robertson2024micro2d]. It was then fine-tuned on the target microstructures of interest: binary carbonate and three-phase SOFC microstructures. Latent representations were encoded as $4\times 16\times 16$ feature maps. The LDM operates on this latent space and was trained for 1000 diffusion steps using a linear noise schedule $\beta\in[1\times 10^{-4},2\times 10^{-2}]$. Full architectural specifications and hyperparameters for both the VAE and LDM are provided in Appendix B.
@@ 3.6 Dataset preparation
For the binary microstructure case, patches were randomly sampled from a $318\times 337$ binary image of a carbonate structure [prodanovic2015digital] to demonstrate 2D-to-3D reconstruction using MicroLad. For the multiphase case, a SOFC anode microstructure [hsu2018mesoscale] was selected, which consists of three phases: pores, nickel (Ni), and yttria-stabilized zirconia (YSZ). In SOFC anodes, each phase plays a distinct role in electrochemical performance: Ni serves as the electron-conducting phase and catalyzes hydrogen oxidation, YSZ is an oxygen-ion conductor that enables ion transport from the electrolyte, and the pore phase provides gas channels for fuel diffusion. A single $950\times 843$ microstructure image was taken from a reference SOFC dataset and manually segmented into the three constituent phases before being used for training. For each case, a total of 300 (binary) and 600 (three-phase) cropped images were randomly sampled, and 8-fold data augmentation was applied to enhance training diversity.
@@ 3.7 Metrics
Relative surface area is a commonly used microstructural descriptor that quantifies the interfacial complexity of each phase, which is particularly important in multiphase materials where phase boundaries influence functional performance (e.g., ion/electron transport pathways in SOFCs) [kench2021generating, hsu2018mesoscale]. In our implementation, surface area is computed by applying Gaussian smoothing to the phase masks and computing their total variation (i.e., the gradient magnitude across the spatial domain) to enable differentiability. The surface area of each phase is then normalized to obtain its relative contribution.
@@
This formulation is sensitive to both phase connectivity and tortuosity [holzer2013redox, zhang2023analysing], making $D_{e}$ an indicator of microstructural functionality. Additional details on the FEM implementation are provided in Section C.
@@ 4 Results
To further validate the reconstruction performance of the proposed MicroLad framework, we compare its accuracy on three-phase SOFC microstructures with two existing methods: MPDD [lee2024denoising, lee2024multi] and SliceGAN [kench2021generating]. As shown in Appendix D, the maximum relative error in terms of two-point correlation for MPDD is 12.44%, and for SliceGAN it is 11.01%. In contrast, MicroLad achieves lower errors when $S_{2}$ guidance is applied, demonstrating strong performance in 2D-to-3D microstructure reconstruction with respect to spatial correlation.
@@ 4.2 Inverse-controlled 2D-to-3D generated microstructure samples
Figure 6 presents the generated 3D microstructure volumes targeting the maximum relative surface area observed in the original dataset for each phase. In contrast to the $V_{f}$-guided generation, the surface area-guided results reveal a tendency for each material phase to form elongated, thin, and interconnected structures that maximize interfacial area with adjacent phases. This behavior also demonstrates that the proposed inverse-controlled generation framework can capture and reflect physically meaningful trends, as the surface area between phases plays a critical role in the electrochemical performance of SOFC anodes by influencing phase boundary reactions [hsu2018mesoscale]. These results highlight MicroLad’s ability to control and analyze phase-specific morphological features and to generate realistic 3D structures suitable for studying interfacial complexity.
@@ 5.1 Distributional shifts and objective trends in latent space
To assess whether generation with the proposed MicroLad framework genuinely expands the microstructure analysis space, we can analyze the property distribution as done in Section 4.2. Additionally, we can examine the data distribution in the latent space of the LDM. To investigate the behavior of the inverse-controlled generation and the distance between the 2D-to-3D generated data (i.e., 2D slices within the 3D volume) and the original 2D data, we examine their latent embeddings using uniform manifold approximation and projection (UMAP) [mcinnes2018umap]. A detailed explanation of the latent embedding procedure and UMAP projection is provided in Appendix F.
@@
Figure 8(h) shows that the trend for increasing the effective diffusivity of the pore phase closely resembles those of $V_{f}^{\text{Pore}}$ and $SA^{\text{Pore}}$. This similarity is expected, as the volume fraction of the pore phase is relatively low compared to the YSZ and Ni phases, and the relative diffusivity of the pore phase in 2D slices is also quite small (with a maximum value around 0.0039). Consequently, achieving a higher $D_{e}^{\text{Pore}}$ effectively requires increasing both the pore volume fraction and surface area. In contrast, Figures 8(i) and 8(j) show more abrupt, threshold-like behavior for $D_{e}^{\text{YSZ}}$ and $D_{e}^{\text{Ni}}$, respectively. In these cases, the effective diffusivity remains low until a certain structural threshold is reached, beyond which the diffusivity rapidly increases. This is physically reasonable: once a given phase becomes sufficiently connected across the domain, forming a continuous pathway from one side to the other, its effective diffusivity improves significantly. This on/off behavior is particularly evident in $D_{e}^{\text{Ni}}$, where a cluster of high-diffusivity samples appears in the lower-left region of the plot in Figure 8(j). Although the diffusivity of YSZ remains lower than that of Ni—primarily due to its granular structure and presence of sintered necks [pecho20153d, hsu2018mesoscale], as opposed to the more continuous network formed by Ni—a similar trend can still be observed in $D_{e}^{\text{YSZ}}$, as shown in Figure 8(i).
@@
To compare the diffusive behavior of each phase in the inverse-controlled generated 3D microstructures with those reconstructed (Recon.) without property guidance, the 3D tortuosity [ghanbarian2013tortuosity, cooper2016taufactor] of both sets of volumes was computed along the three Cartesian axes (x, y, and z), as shown in Table 2. As observed, the reconstructed 3D volumes generally exhibit higher tortuosity (i.e., lower diffusivity), while the inverse-controlled samples show reduced tortuosity as the target $D_{e}$ is optimized for each material phase of interest. This observation suggests that enforcing target diffusivity at the slice level during 3D microstructure generation can effectively promote globally connected diffusive networks within the volume. Future work should explore more advanced control strategies, such as directionally weighted SDS loss or hierarchical optimization across multi-scale representations, to further enhance the spatial tailoring of 3D transport properties. More sophisticated control over 3D material behavior also warrants further investigation.